According to the Can Manufacturers Institute, the energy used to make an aluminum can from recycled aluminum is 5% of the energy used to make an aluminum can from virgin ore. In a typical year, 1.7 billion pounds of aluminum cans are recycled.

Part A

How much energy is thermally transferred to get this mass of aluminum from 20 ∘C to its melting point, 660 ∘C?
Answer #1

we use the relation, Q =mc\DeltaT

m =1.7billion pounds = 7.7*108kg

c = specific heat capacity = 0.90 J oC-1 g-1

so, Q = 7.7*1011 * 0.90 * (660-20) = 4.4*1014J